OpenStudy (anonymous):

the ratio of the number of people who own a smartphone to the number of people who own a flip phone is 4:3.If 500 more people own a smartphone then flip phone,how many people own each type of phone. i give a medal if u help me

OpenStudy (gibbs):

10

OpenStudy (anonymous):

4(smartphone):3(flip phone) = 500(smartphone):x(flip phone). Divide 500 by 4 which equals 125. Then multiply 125 by 3 to get 500:375. That means for 500 people using a smartphone there are 375 people using a flip phone.

OpenStudy (anonymous):

so what should i write on my papaer

OpenStudy (anonymous):

@Dscdago if u tell what to write and help me i will give u a medal and a fan

OpenStudy (anonymous):

Since a ratio is basically a fraction. \[\frac{ 4 }{ 3 }=\frac{ 500 }{ x }\] divide 500 by 4 to find by what number you multiply 4 to get 500. You multiply that same number by 3 to get x. 500/4 = 125. \[\frac{ 4(125) }{3(125) }=\frac{ 500 }{ 375 }\] So for every 500 smartphone user, there is 375 people using a flip phone.

OpenStudy (anonymous):

i gave u a medal and should i write that in my notebook

OpenStudy (anonymous):

Yes. You can use this method for almost every ratio problem.

OpenStudy (anonymous):

thanks i am fan of u i gave u medal when ever i ask u for help would u help me please

OpenStudy (anonymous):

Sure whenever I'm available.. I may be helping other people or I might be working on my own work.

OpenStudy (anonymous):

bye cute pie i am a girl